OpenStudy (anonymous):

integrate 1/(x^2) sinh (1/x) dx

OpenStudy (solomonzelman):

u=1/x

OpenStudy (solomonzelman):

smooth

OpenStudy (anonymous):

then 1/x^2 will become u^2?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\color{red}{\frac{1}{x^2}}\color{blue}{\sinh\left(\frac{1}{x}\right)}\color{red}{dx} }\) \(\color{#0000ff }{ \displaystyle u=1/x }\) \(\color{#ff0000 }{ \displaystyle du=-1/x^2{\tiny~}dx\quad \Longrightarrow \quad -du=1/x^2{\tiny~}dx }\)

OpenStudy (solomonzelman):

substitute accordingly... you are probably just too tired working these out :)

OpenStudy (solomonzelman):

the (1/x²)dx goes away and is replaced by -du. And 1/x inside sinh is replied by u.

OpenStudy (anonymous):

ill do it again,

OpenStudy (solomonzelman):

(I shouldn't have coloered "sinh" in blue) ...\ sure :)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int f'(x)\cdot {\bf G}\left\{f(x)\right\} dx}\) \(\color{#000fc0 }{ \displaystyle u=f(x)}\) \(\color{#000fc0 }{ \displaystyle du=f'(x){\tiny~}dx}\) \(\color{#000000 }{ \displaystyle \int {\bf G}\left\{u\right\} du}\)

OpenStudy (solomonzelman):

that is typically the setup ... your probably isn't different ... (I've seen you do harder integrals before)

OpenStudy (anonymous):

ok, i got it

OpenStudy (anonymous):

yayy

OpenStudy (solomonzelman):

Alright; nice to hear that!

OpenStudy (anonymous):

thanks